# An algorithmic version of the Hajnal–Szemerédi theorem

### v2

@@ 6.5. Auxiliary lemmas
Next, assume $|B^{1}\setminus V(K)|-(r-s)(n/r-|K^{1}|)>0$. In this case, recall that $K$ contains a copy of $K_{r}$, denoted by $H$, such that $|V(H)\cap B^{1}|=r-s+1=3$. Now if $G[B^{3}\cup(V(H)\cap B^{1})]$ is connected, then by Proposition 6.17, let $H_{1}$ be a copy of $K_{r}$ that consists of $V(H)\setminus B^{1}$ and a triangle from $B^{*}$. Then either $G[(B^{3}\cup V(H))\setminus V(H_{1})]$ is connected or it has two even components, and in either case, it has a perfect matching. Then we can find a $K_{r}$-factor of $G[A_{1}^{3}\cup\cdots\cup A_{s}^{3}\cup(B^{3}\cup V(H))\setminus V(H_{1})]$ by Lemma 6.3 similarly as above, which together with $((K\cup K^{b})\setminus\{H\})\cup\{H_{1}\}$ yields a $K_{r}$-factor of $G$. The remaining case is that $G[B^{3}\cup(V(H)\cap B^{1})]$ is disconnected. Without loss of generality, assume that $V(H)\cap B^{1}$ has neighbors only in $B^{*}$. Then by Proposition 6.17, let $H_{2}$ be a copy of $K_{r}$ that consists of $V(H)\setminus B^{1}$ and a triangle from $B^{**}$. Note that $G[(B^{3}\cup V(H))\setminus V(H_{2})]$ has two even components and we can find a $K_{r}$-factor of $G$ by taking the union of the $K_{r}$-factor returned by Lemma 6.3 and $((K\cup K^{b})\setminus\{H\})\cup\{H_{2}\}$. ∎
@@
If $r-s=2$ and $B^{1}_{b}\neq\emptyset$, then we can find a $K_{r}$-factor of $G$ in time $O(n^{r})$.
@@
Next we cover the $x-1$ bad vertices except $v$ by a small $K_{r}$-tiling. Indeed, we use Lemma 6.14 repeatedly to find a $K_{r}$-tiling $F$ whose members are all balanced copies of $K_{r}$ and contain all these bad vertices. Let $\tilde{B}$ be the remaining of vertices of $B^{1}$, excluding $v$. Clearly, $|\tilde{B}|$ is odd and $\tilde{B}\subseteq B_{g}^{1}$. Note that $d_{\tilde{B}}(v)\geq 0.9n/r$. If $\tilde{B}$ is connected, then there exists a vertex $y\in\tilde{B}\cap N(v)$ such that $\tilde{B}\setminus\{y\}$ is still connected. Indeed, if such $y$ does not exist, then all neighbors of $v$ are cut-vertices. In particular, $\tilde{B}$ has two cut-vertices, which leaves a 2-connected component $Z$ of size at most $2n/(3r)$. Then any vertex in $Z$ has degree at most $2n/(3r)+2<0.99n/r$ in $\tilde{B}$, a contradiction. Then we can find a balanced copy $T$ of $K_{r}$ containing the edge $vy$, which is possible by (P2)(P5). On the other hand, if $\tilde{B}$ is disconnected, then note that it has an even component $X$ and an odd component $Y$. Recalling that $\delta(G[B^{3}])\geq n/r-6s\mu^{3/5}n-c$ and $|F|\leq r$, we infer that $|Y|\geq n/r-6s\mu^{3/5}n-c-2r$. By (6.1), we have $d_{Y}(v)>0$. Then similarly we pick a balanced copy $T$ of $K_{r}$ containing the edge $vy$, where $y\in Y$ by  (P2)(P5). Therefore, in both cases, we have that $\tilde{B}\setminus V(T)=\tilde{B}\setminus\{y\}$ has a perfect matching.
@@
It remains to consider the case $r-s=2$, $I=[s+1]$ and $B^{1}_{b}=\emptyset$. Since $I=I^{*}=[s+1]$, $K^{1}$ is indeed $[s+1]$-admissible and $K=K^{1}$. Moreover, since $B^{1}_{b}=\emptyset$, we have $K^{b}=\emptyset$.
@@
If $r-s=2$, $I=[s+1]$, $B^{1}_{b}=\emptyset$ and $B^{1}$ is not disconnected with two odd components, then we can find a $K_{r}$-factor of $G$ in time $O(n^{r})$.
@@
Let $B_{1}^{*}:=\{v\in B^{1}:d_{B^{*}}(v)\geq 0.1|B^{1}|\}$ and $B_{1}^{**}:=\{v\in B^{1}:d_{B^{**}}(v)\geq 0.1|B^{1}|\}$. Note that we can find $B_{1}^{*}$ and $B^{**}$ in time $O(n^{2})$. By the minimum degree condition of $B^{1}$, we know $B^{1}=B_{1}^{*}\cup B_{1}^{**}$. Moreover, $B^{*}\subseteq B_{1}^{*}$ and $B^{**}\subseteq B_{1}^{**}$.
@@
Therefore, the proof of (1) in the lemma is completed, and it remains to consider the case that $B^{1}$ is disconnected with two odd components. We need to show that $G$ has a $K_{r}$-factor if and only if $G$ has a strong admissible $K_{r}$-tiling. Moreover, the existence of such $K_{r}$-tiling in $G$ can be checked by an algorithm that runs in time $n^{O(1)}2^{O(c^{{r}^{r+2}})}$, and if $G$ has a $K_{r}$-factor, it also outputs a $K_{r}$-factor.
@@
To check whether $G$ has a strong admissible $K_{r}$-tiling, we fix a copy $H$ of $K_{r}$ with an odd number of vertices of $B^{1}$ and use Lemma 5.2 to search for a $K_{r}$-tiling $F$ in $G\setminus V(H)$ such that $F\cup\{H\}$ is $[s+1]$-admissible and $|F\cup\{H\}|\leq(3t_{[s+1]})^{r}+2\leq(r^{2}c/0.11)^{r^{r+2}}+2$. By similar arguments in the proof of Lemma 6.12, the color-coding search takes time $n^{O(1)}2^{O(c^{{r}^{r+2}})}$. Since there are at most $n^{r}$ such $H$, the running time of this process is $O(n^{r})\cdot n^{O(1)}2^{O(c^{{r}^{r+2}})}=n^{O(1)}2^{O(c^{{r}^{r+2}})}$.
@@
Now suppose $G$ has no strong admissible $K_{r}$-tiling and we need to show that $G$ has no $K_{r}$-factor. We may assume that $G$ has an $[s+1]$-admissible $K_{r}$-tiling $F$ of size at most $(3t_{[s+1]})^{r}$ – otherwise, by Lemma 1.7, $G$ has no $K_{r}$-factor. Since $F$ is not a strong admissible $K_{r}$-tiling, every copy of $K_{r}$ in $F$ has an even number of vertices of $B^{1}$. Moreover, we claim that there is no edge in $A_{i}^{1}\setminus V(F)$ for each $i\in[s]$. Suppose not and without loss of generality, $A_{1}^{1}\setminus V(F)$ has an edge. Then we can find two disjoint copies of $K_{r}$ with index vector $(2,1,\ldots,1,1)$ and $(0,1,\ldots,1,3)$, respectively. The union of $F$ and these two copies of $K_{r}$ is a strong admissible $K_{r}$-tiling, a contradiction. This implies that $G-F$ contains $s$ independent sets $A_{i}^{1}\setminus V(F)$ for $i\in[s]$ and $B^{1}\setminus V(F)$, which consists of two odd components. Then note that $G-F$ does not induce a $K_{r}$-factor. Indeed, no copy of $K_{r}$ in $G-F$ contains more than $r-2$ vertices from $\bigcup_{i\in[r-2]}A_{i}^{1}\setminus V(F)$, and since it is balanced, a $K_{r}$-factor of $G-F$ (if it exists) must only contain balanced copies of $K_{r}$. This would induce a perfect matching in $B^{1}\setminus V(F)$, which does not exist. Since this holds for any such $F$ of size at most $(3t_{[s+1]})^{r}$, by Lemma 1.7, $G$ has no $K_{r}$-factor.
@@ 6.7. When $s=r$
Given an $n$-vertex graph $G$ with $\delta(G)\geq(1-1/r)n-c$, let $(A_{1}^{1},\dots,A_{s}^{1},B^{1})$ be the partition defined above in this section and suppose $s=r$. Let $(A_{1}^{2},\ldots,A_{r}^{2})$ and $I^{*}\subseteq[r]$ be defined as above. Then $G$ has a $K_{r}$-factor if and only if $G$ contains an $I^{*}$-admissible $K_{r}$-tiling. Moreover, the existence of such $K_{r}$-tiling in $G$ can be checked by an algorithm that runs in time $n^{O(1)}2^{O(c^{{r}^{r+2}})}$, and if $G$ has a $K_{r}$-factor, it also outputs a $K_{r}$-factor.
@@
We start with applying Lemma 6.12 to search for an $I^{*}$-admissible $K_{r}$-tiling $K_{1}$ in time $n^{O(1)}2^{O(c^{r^{r+2}})}$. If it does not output an $I^{*}$-admissible $K_{r}$-tiling, then by Lemma 1.7, $G$ has no $K_{r}$-factor (and the partition $(A_{1}^{2},\ldots,A_{r}^{2})$ is a certificate). So we may assume that it outputs an $I^{*}$-admissible $K_{r}$-tiling $K_{1}$. Then we apply Lemma 6.13 to find an $[r]$-admissible $K_{r}$-tiling $K\supseteq K^{1}$ of size at most $r^{2}\mu^{2/3}n$ such that $K^{1}\subseteq K$ and $V(G)\setminus V(K)$ has a partition $A_{1}^{3},\ldots,A_{r}^{3}$ with $|A_{1}^{3}|=\cdots=|A_{r}^{3}|\geq n/r-r^{2}\mu^{2/3}n$, $A_{i}^{3}\subseteq A_{i}^{2}$ for each $i\in[r]$.

### v3

@@ 6.5. Auxiliary lemmas
Next, assume $|B^{1}\setminus V(K^{1})|-(r-s)(n/r-|K^{1}|)>0$. In this case, recall that $K$ contains a copy of $K_{r}$, denoted by $H$, such that $|V(H)\cap B^{1}|=r-s+1=3$. Now if $G[B^{3}\cup(V(H)\cap B^{1})]$ is connected, then by Proposition 6.17, let $H_{1}$ be a copy of $K_{r}$ that consists of $V(H)\setminus B^{1}$ and a triangle from $B^{*}$. Then either $G[(B^{3}\cup V(H))\setminus V(H_{1})]$ is connected or it has two even components, and in either case, it has a perfect matching. Then we can find a $K_{r}$-factor of $G[A_{1}^{3}\cup\cdots\cup A_{s}^{3}\cup(B^{3}\cup V(H))\setminus V(H_{1})]$ by Lemma 6.3 similarly as above, which together with $((K\cup K^{b})\setminus\{H\})\cup\{H_{1}\}$ yields a $K_{r}$-factor of $G$. The remaining case is that $G[B^{3}\cup(V(H)\cap B^{1})]$ is disconnected. Without loss of generality, assume that $V(H)\cap B^{1}$ has neighbors only in $B^{*}$. Then by Proposition 6.17, let $H_{2}$ be a copy of $K_{r}$ that consists of $V(H)\setminus B^{1}$ and a triangle from $B^{**}$. Note that $G[(B^{3}\cup V(H))\setminus V(H_{2})]$ has two even components and we can find a $K_{r}$-factor of $G$ by taking the union of the $K_{r}$-factor returned by Lemma 6.3 and $((K\cup K^{b})\setminus\{H\})\cup\{H_{2}\}$. ∎
@@
If rs=2r-s=2 and $B^{1}_{b}\neq\emptyset$, then we can find a $K_{r}$-factor of $G$ in time $O(n^{r+1})$.
@@
[…]
@@
It remains to consider the case rs=2r-s=2, $I=[s+1]$ and $B^{1}_{b}=\emptyset$. Since $I=I^{*}=[s+1]$, $K^{1}$ is indeed $[s+1]$-admissible and $K=K^{1}$. Moreover, since $B^{1}_{b}=\emptyset$, we have $K^{b}=\emptyset$.
@@
If rs=2r-s=2, $I=[s+1]$, $B^{1}_{b}=\emptyset$ and $B^{1}$ is not disconnected with two odd components, then we can find a $K_{r}$-factor of $G$ in time $O(n^{r+1})$.
@@
Let $B_{1}^{*}:=\{v\in B^{1}:d_{B^{*}}(v)\geq 0.1|B^{1}|\}$ and $B_{1}^{**}:=\{v\in B^{1}:d_{B^{**}}(v)\geq 0.1|B^{1}|\}$. Note that we can find $B_{1}^{*}$ and $B_{1}^{**}$ in time $O(n^{2})$. By the minimum degree condition of $B^{1}$, we know $B^{1}=B_{1}^{*}\cup B_{1}^{**}$. Moreover, $B^{*}\subseteq B_{1}^{*}$ and $B^{**}\subseteq B_{1}^{**}$.
@@
Therefore, the proof of (1) in the lemma is completed, and it remains to consider the case that $B^{1}$ is disconnected with two odd components. We need to show that $G$ has a $K_{r}$-factor if and only if $G$ has a strong admissible $K_{r}$-tiling. Moreover, the existence of such $K_{r}$-tiling in $G$ can be checked by an algorithm that runs in time $n^{O(1)}2^{O(c^{{r}^{r+2}})}$ for $r\geq 4$ (resp. $n^{O(1)}2^{O(c)}$ for $r=3$), and if $G$ has a $K_{r}$-factor, it also outputs a $K_{r}$-factor.
@@
To check whether $G$ has a strong admissible $K_{r}$-tiling, we fix a copy $H$ of $K_{r}$ with an odd number of vertices of $B^{1}$ and use Lemma 5.4 to search for a $K_{r}$-tiling $F$ in $G\setminus V(H)$ such that $F\cup\{H\}$ is $[s+1]$-admissible and $|F\cup\{H\}|\leq g(t_{[s+1]})+2$. By similar arguments in the proof of Lemma 6.12, the color-coding search takes time $n^{O(1)}2^{O(g(t_{[s+1]}))}$. Since there are at most $n^{r}$ such $H$, the running time of this process is $O(n^{r})\cdot n^{O(1)}2^{O(g(t_{[s+1]}))}=n^{O(1)}2^{O(g(t_{[s+1]}))}$, which is $n^{O(1)}2^{O(c^{r^{r+2}})}$ for $r\geq 4$ and $n^{O(1)}2^{O(c)}$ for $r=3$.
@@
Now suppose $G$ has no strong admissible $K_{r}$-tiling and we need to show that $G$ has no $K_{r}$-factor. We may assume that $G$ has an $[s+1]$-admissible $K_{r}$-tiling $F$ of size at most $g(t_{[s+1]})$ – otherwise, by Lemma 1.8, $G$ has no $K_{r}$-factor. Since $F$ is not a strong admissible $K_{r}$-tiling, every copy of $K_{r}$ in $F$ has an even number of vertices of $B^{1}$. Moreover, we claim that there is no edge in $A_{i}^{1}\setminus V(F)$ for each $i\in[s]$. Suppose not and without loss of generality, $A_{1}^{1}\setminus V(F)$ has an edge. Then we can find two disjoint copies of $K_{r}$ with index vector $(2,1,\ldots,1,1)$ and $(0,1,\ldots,1,3)$, respectively. The union of $F$ and these two copies of $K_{r}$ is a strong admissible $K_{r}$-tiling, a contradiction. This implies that GFG-F contains $s$ independent sets $A_{i}^{1}\setminus V(F)$ for $i\in[s]$ and $B^{1}\setminus V(F)$, which consists of two odd components. Then note that GFG-F does not induce a $K_{r}$-factor. Indeed, no copy of $K_{r}$ in GFG-F contains more than r2r-2 vertices from $\bigcup_{i\in[r-2]}A_{i}^{1}\setminus V(F)$, and since it is balanced, a $K_{r}$-factor of GFG-F (if it exists) must only contain balanced copies of $K_{r}$. This would induce a perfect matching in $B^{1}\setminus V(F)$, which does not exist. Since this holds for any such $F$ of size at most $g(t_{[s+1]})$, by Lemma 1.8, $G$ has no $K_{r}$-factor.
@@ 6.7. When $s=r$
Given an $n$-vertex graph $G$ with $\delta(G)\geq(1-1/r)n-c$, let $(A_{1}^{1},\dots,A_{s}^{1},B^{1})$ be the partition defined above in this section and suppose $s=r$. Let $(A_{1}^{2},\ldots,A_{r}^{2})$ and $I^{*}\subseteq[r]$ be defined as above. Then $G$ has a $K_{r}$-factor if and only if $G$ contains an $I^{*}$-admissible $K_{r}$-tiling. Moreover, the existence of such $K_{r}$-tiling in $G$ can be checked by an algorithm that runs in time $n^{O(1)}2^{O(c^{{r}^{r+2}})}$ for $r\geq 4$ (resp. $n^{O(1)}2^{O(c)}$ for $r=3$), and if $G$ has a $K_{r}$-factor, it also outputs a $K_{r}$-factor.
@@
We start with applying Lemma 6.12 to search for an $I^{*}$-admissible $K_{r}$-tiling $K_{1}$ in time $n^{O(1)}2^{O(c^{{r}^{r+2}})}$ for $r\geq 4$ (resp. $n^{O(1)}2^{O(c)}$ for $r=3$). If it does not output an $I^{*}$-admissible $K_{r}$-tiling, then by Lemma 1.8, $G$ has no $K_{r}$-factor (and the partition $(A_{1}^{2},\ldots,A_{r}^{2})$ is a certificate). So we may assume that it outputs an $I^{*}$-admissible $K_{r}$-tiling $K_{1}$. Then we apply Lemma 6.13 to find an $[r]$-admissible $K_{r}$-tiling $K\supseteq K^{1}$ of size at most $r^{2}\mu^{2/3}n$ such that $K^{1}\subseteq K$ and $V(G)\setminus V(K)$ has a partition $A_{1}^{3},\ldots,A_{r}^{3}$ with $|A_{1}^{3}|=\cdots=|A_{r}^{3}|\geq n/r-r^{2}\mu^{2/3}n$, $A_{i}^{3}\subseteq A_{i}^{2}$ for each $i\in[r]$.